Calculate the hardness of water sample whose 10 ml required 20 ml of EDTA solution. 20 ml of CaCl2 solution whose strength is equivalent to 1.5 g of CaCO3 per litre, required 30 ml of EDTA solution

Grade:12th pass

4 Answers

Pawan Prajapati
3 years ago
Answer ::

The standard water contains 1.5 g of CaCO3/l.
1000 ml of standard hard water (SHW), therefore, contain 15,00 mg of CaCO3.
1 ml of standard water = 15,00/1000 = 1.5 mg of CaCO3 equivalent.
10 ml of EDTA solution = 10 ml of SHW contains
=> 10* 1.5 = 15 mg of CaCO3 equivalent
1ml o EDTA Solution = 15 / 10 = 1.5 mg of CaCo3 equivalent
Calculation of hardness of water:
100 ml of water 30 ml of EDTA solution
30 * 1.5 = 45 mg of CaCO3 equivalent
1000 ml of water = 450 mg of CaCO3 equivalent
Hardness of water = 450 mg / l

Thanks
Tamilarasan
3 years ago
1ml of cacl2=1.5  mg of caco3
30 ml of cacl2 =20 mg of caco3
20 ml of cacl2 = 20*1.5 =30 mg of caco3
1 ml of cacl2 =30/30 =1 mg of caco3
10 ml of sample water = 20 ml of edta
Total hardness,
Hardness=( 20/100)*1000
         =2000 ppm
irshad hussain
2 years ago

he standard water contains 1.5 g of CaCO3/l.

1000 ml of standard hard water (SHW), therefore, contain 15,00 mg of CaCO3.

1 ml of standard water = 15,00/1000 = 1.5 mg of CaCO3 equivalent.

10 ml of EDTA solution = 10 ml of SHW contains

=> 10* 1.5 = 15 mg of CaCO3 equivalent

1ml o EDTA Solution = 15 / 10 = 1.5 mg of CaCo3 equivalent

Calculation of hardness of water:

100 ml of water 30 ml of EDTA solution

30 * 1.5 = 45 mg of CaCO3 equivalent

1000 ml of water = 450 mg of CaCO3 equivalent

Hardness of water = 450 mg / l

irshad hussain
2 years ago

the standard water contains 1.5 g of CaCO3/l.

1000 ml of standard hard water (SHW), therefore, contain 15,00 mg of CaCO3.

1 ml of standard water = 15,00/1000 = 1.5 mg of CaCO3 equivalent.

10 ml of EDTA solution = 10 ml of SHW contains

=> 10* 1.5 = 15 mg of CaCO3 equivalent

1ml o EDTA Solution = 15 / 10 = 1.5 mg of CaCo3 equivalent

Calculation of hardness of water:

100 ml of water 30 ml of EDTA solution

30 * 1.5 = 45 mg of CaCO3 equivalent

1000 ml of water = 450 mg of CaCO3 equivalent

Hardness of water = 450 mg / l